jaime has available a 10% alcohol solution and a 60% alcohol solution. Find how many liters of each solution he should mix to make 50 liters of a 40% alcohol solution
Answer by stanbon(75887)

Let amount of 10% solution be x ; 0.10x is alcohol
Amount of 60% solution must be "50-x" ; 0.60(50-x) is alcohol
EQUATION:
alcohol + alcohol = .40(50)= 20 liters of alcohol
0.10x+0.6(50-x)= 20
Multiply thru by 10 to get:
x+6(50-x)=200
x+300-6x=200
-5x=-100
x=20 (liters of 10% solution)
50-x=30 (liters of 60% solution)
Cheers,
Stan H.